the sum of two numbers is 48 the first is 8 less than the second. find the numbers
Answer by poweezo(4)

x(x-8)=48
x^2-8x=48
x^2-8x-48=0
(x-12)(x+4)=0
x-12=0 or x+4=0
x=12 or x=-4
The first number is 4 and the second is 12